# LCI: a Lightweight Communication Interface for Efficient Asynchronous Multithreaded Communication

### v1

@@ 4.5. Implementation Note
LCI is implemented as a C++17 library with the CMake build system. It is also available as a Spack package. LCI supports four bootstrapping backends: PMI1, PMI2, PMIx, and MPI. It has been tested on Infiniband, RoCE, Slingshot-11, and Ethernet networks. It is fully open-sourced with the MIT license.
@@ 5.1. Experimental Setup
We evaluate LCI on SDSC Expanse and NCSA Delta. Table 2 shows their configuration. Expanse uses InfiniBand, which is one of the most widely used interconnects for HPC clusters and accounts for 61% of the Top500 systems. Delta uses Slingshot-11, which is increasingly popular and used on 7 of the top-10 systems. [^1] Statistics are based on the TOP500 List published in Nov. 2024. All experiments are conducted at least six times. The figures show the average and standard deviation.
@@ 5.2. Micro-benchmarks
In asynchronous multithreaded applications, message rate and bandwidth are more critical than latency due to communication overlapping and nonblocking execution. Therefore, we use these two metrics to compare LCI with standard MPI, MPICH with the VCI extension, and GASNet-EX. Our micro-benchmarks run on two nodes with two basic modes. The process-based mode uses one process on each core, while the thread-based setting uses one process on each node with one thread per core. Each process/thread has a peer process/thread on the other node, and it performs ping-pongs with the peer. Existing multithreaded applications can either share a global set of communication resources or, if the application logic and underlying communication library permit, allocate dedicated resources for each thread. Therefore, the thread-based mode is further divided into two sub-modes according to the resource-sharing pattern: (a) in the dedicated resource mode, each thread allocates its communication resources; (b) In the shared resource mode, all threads share a global set of communication resources. The dedicated resource mode is implemented with MPICH VCIs and LCI devices. Cray-MPICH and GASNet-EX do not support this mode. We also set mpi_assert_no_any_tag and mpi_assert_allow_overtaking to true and configure MPIR_CVAR_CH4_GLOBAL_PROGRESS to 0 to minimize the thread contention on VCIs.
To ensure uniformity across different communication libraries, we build a simple layer (the Lightweight Communication Wrapper, or LCW) on top of LCI, MPI, and GASNet-EX and use it to write the microbenchmarks. The microbenchmarks, along with the LCW layer, are open-sourced [^2] https://github.com/¡anonymous¿/lcw. LCW implements simple nonblocking active messages and send-receive primitives. For MPI, it uses MPI_Isend/MPI_Irecv for send-receive and MPI_Isend/pre-posted MPI_Irecv for active messages. For GASNet-EX, it uses gex_AM_RequestMedium for active messages and does not support send-receive due to implementation complexity. We show the active message results in the message rate microbenchmark and the send-receive results in the bandwidth microbenchmark.
5.2.1. Single-threaded Performance
@@
Figure 2. Process-based message rate micro-benchmark. We use one process per core and one thread per process.
Figure 2 shows the single-threaded message rate results. We fixed the message size to 8 bytes and increased the process number from 1 to 128 per node. Each process/thread runs 100k iterations. We report the uni-directional message rate. LCI achieves performance comparable to the other communication libraries. Figures for the single-threaded bandwidth results are omitted due to page limit, but the results are similar.
@@ 5.2.2. Multithreaded Performance
Figure 3. Thread-based message rate micro-benchmark. We use one process per node and one thread per core. Dedicated resources uses one LCI device/MPICH VCI per thread. Shared resources uses one set of resources for the entire process.
Figure 3 shows the multithreaded message rate results. We fixed the message size to 8 bytes and increased the thread number from 1 to 128 per node. LCI achieves significant speedups in multithreaded performance on both platforms (sometimes more than 10x). In particular, multithreaded LCI with dedicated devices achieves even slightly better performance than multi-process LCI (around 15% at full scale). The MPICH’s VCI extension greatly helps multithreaded performance, but the overall performance is still suboptimal. GASNet-EX shows good multithreaded performance in the shared resource mode, but its lack of resource-replication support weakens its competencies if the application wants to use more resources.
Even though we do not directly evaluate UCX and Libfabric due to the difficulty of bootstrapping and the complexity of their APIs, the MPICH results on Expanse (particularly Figure 3(a)) give hints on their multithreaded performance. UCX is generally faster than libfabric on InfiniBand, but its performance degrades sharply when there are more than 16 threads. Libfabric shows good scaling results with dedicated resources at the cost of absolute performance numbers. LCI achieves the best of both worlds by directly building on the lowest-level public API, libibverbs. UCX does not support Slingshot-11, so its results on Delta are unavailable. MPICH does not support more than 64 VCIs, so some data points are missing.
@@
Figure 4. Thread-based bandwidth micro-benchmark. We use one process per node and one thread per core. Dedicated resources uses one LCI device/MPICH VCI per thread. Shared resources uses one set of resources for the entire process.
@@ 5.2.3. Individual Resources
LCI communications involve operations on a variety of resources. Each resource is optimized for threading efficiency, and users can explicitly allocate multiple replicas of them. Our next microbenchmark evaluates the threading efficiency of three major LCI resources: completion queue, matching engine, and packet pool. All microbenchmarks run on a single node on Delta with different thread numbers. All threads perform 100k of key resource methods that are used in the communication critical path (a pair of completion queue push/pop, matching engine inserts, or packet pool get/put). Figure 5 shows the results. As we can see, the packet pool and matching engine scales well with thread number, achieving 800 Mops (Million operations per second) or 260 Mops with 128 threads. As a reference, our ping-pong microbenchmark achieves at most 22 Million Messages per second (Figure 3(a)). This means allocating one instance of each resource per process is sufficient. The completion queue achieves 18 Mops with 128 threads, which means applications aiming for higher throughput may need to allocate more completion queues per process. The completion queue throughput is primarily constrained by how fast threads can perform the atomic fetch-and-add operation on a shared variable. Our message rate microbenchmark shown above uses one completion queue per thread.
@@ 5.3. K-mer Counting
In the k-mer counting stage, HipMer traverses the dataset twice. The first traversal inserts the k-mers into a two-layer Bloom filter. A Bloom filter is a space-efficient data structure that tests whether an element is in a set with a small false positive rate. The second traversal then consults the Bloom filter and inserts those with more than one occurrence into a hashmap. The hashmap maintains the actual count of the k-mers, while the two-layer Bloom filter is used to reduce the memory footprint of the hashtable by filtering out those occurring only once (which are likely erroneous).
HipMer is written in UPC++ with only one thread per process. Each k-mer is statically mapped to a process using a hash function. Each process reads part of the dataset and sends the k-mers to the mapped processes via UPC++ RPCs. It further employs an aggregation buffer per target process to reduce communication overhead.
We implement a multithreaded version of the HipMer k-mer counting stage. The new implementation is also based on the RPC abstraction and aggregation, with libcuckoo hashtable(Li et al., 2014) and a hand-written atomic-based Bloom filter. It supports two network backends, LCI and GASNet-EX, primarily leveraging their active message primitives. The LCI backend shares many similarities with the one described in Section 3.2. Compared to the single-threaded implementation, multithreading reduces the number of aggregation targets by a factor of N, where N is the thread number per process. All threads can serve the incoming RPCs, resulting in improved load balance.
@@
We run the k-mer counting mini-app with the human chr14 dataset (7.75GB). It contains 37 million reads and 1.8 billion k-mers (with the k-mer length $k=51$). We run the multithreaded implementation with 2 processes per node to avoid the inter-socket overheads. The aggregation buffer size is set to be 8KB per destination. The total aggregation buffer size is always smaller than its HipMer counterpart due to the reduced destination number. All threads run the application logic and periodically progress the network backend (the all-worker setup). This is the best setup for LCI on both platforms. However, when running GASNet-EX on Delta, the all-worker setup results in devastating performance (over 20x worse than LCI). Therefore, we add an additional dedicated progress setup for GASNet-EX: use 63 threads for application logic and one thread for network progress. We report the better of the two setups for GASNet-EX (all-worker setup on Expanse and dedicated progress setup on Delta).
Fig.6 shows the strong scaling results of the mini-app on Expanse and Delta from 1 node (2 processes/128 cores) to 32 nodes (64 processes/4096 cores). Our multithreaded implementation outperforms the single-threaded reference implementation by up to 60% on Expanse (8 nodes) and 40% on Delta (4 nodes), at which point the reference implementation suffers from severe load imbalance problems across 512/1024 processes. In addition, the LCI backend outperforms its GASNet-EX counterpart by 35% on Expanse (16 nodes) and 55% on Delta (16 nodes). Although not shown here, we also tried larger aggregation buffer sizes (up to 64KB), which resulted in slightly smaller gaps between GASNet-EX and LCI due to less frequent communication, but lower overall performance due to worse load balance.
HipMer evaluation stops at 8/4 nodes because UPC++ takes too long to bootstrap for larger process counts. Investigation shows it was due to the slow PMI2 fence operation. It reflects a common challenge of the one-process-per-core running mode on modern HPC systems with high intra-node parallelism.
@@ 5.4. HPX and Octo-Tiger
The increasingly complicated architectures and dynamic scientific computing algorithms have attracted growing interest in the Asynchronous Many-Task (AMT) programming model (Kaiser et al., 2023; Bauer et al., 2 11; Kale and Krishnan, 0 01; Augonnet et al., 2009, 2024). With AMTs, users express their application logic as a set of fine-grained tasks and task dependencies. The runtime then schedules these tasks on available resources according to task dependencies and data locality. Compared to the traditional bulk-synchronous parallel (BSP) model, AMTs can potentially achieve better load balance, portability, and communication overlapping, with lower user programming complexity. However, previous works have shown that existing communication libraries usually do not support AMTs’ communication needs most efficiently, as their communications are heavily multithreaded and asynchronous (Castillo et al., 2019; Swartvagher, 2022; Yan et al., 2025).
(Yan et al., 2023) has integrated a previous C version of LCI into HPX (Kaiser et al., 2023), an established AMT runtime that fully complies with the C++ Standard APIs and extends them to the distributed case. In this work, we upgrade the LCI support inside HPX to the latest C++ version and evaluate its performance with an astrophysics application, Octo-Tiger (Marcello et al., 2021), which simulates the evolution of stellar systems based on adaptive octo-trees and fast multipole methods. Octo-Tiger is built on top of HPX for fully asynchronous execution and communication overlapping. We use the ”rotating star” scenarios and report the timestamp per step.
@@
Figure 7 shows the results. For mpix, we use the MPICH VCI extension. Preliminary experiments have shown that MPICH with VCI extensions performs better with the libfabric backend than the UCX backend. Therefore, we use the MPICH libfabric backend here. Results reported here use the optimal VCI number for mpix and the optimal device number for lci. We also use replicated request pools for mpix to reduce thread contention on completion polling. On Expanse, LCI outperforms mpi (standard MPI) by 30% and mpix by 10%. On Delta, LCI outperforms mpi by 3x and mpix by 35%. In addition, mpix needs 8 VCIs on both platforms to reach the optimal performance, while lci only needs 1 device on Expanse and 2 devices on Delta. This shows that LCI has better intra-resource threading efficiency compared to MPICH, thanks to its thread-efficient runtime design.
Related work
A previous version of LCI has been integrated into PaRSEC and showed favorable performance over its MPI backend (Mor et al., 2023). LCI has also been integrated into HPX releases and used by other projects. (Strack et al., 2024) implements a 2D FFT mini-app with HPX and reports that the LCI backend outperforms the MPI counterpart and the reference FFTW implementation by 5x. (Daiß et al., 2024) scales Octo-Tiger to 1700+ GPU nodes (around 7000 GPUs/processes) on Perlmutter and achieves 1.7x speedup compared to Cray MPICH at full scale. We are unable to conduct similar scaling experiments in this paper due to computation resource limitations, but we expect the latest version of LCI to have similar scalability.
@@ 6. Conclusion and Future Work
LCI is still under active development. There are several areas for future improvements. We list some of the most important ones below.
Collective Communication: The existing LCI API focuses on point-to-point communication primitives, as they are more commonly seen in asynchronous multithreaded applications and are also the basic building blocks for collective communication. LCI offers a few basic collective communication primitives, including dissemination-based barrier and tree-based broadcast/reduce. Users can also combine LCI with other communication libraries, such as MPI or *CCL, for better collective communication performance. Identifying the right interface and efficient design of non-bulk-synchronous collective communication for asynchronous multithreaded applications is an important future work.
GPU Communication: The existing LCI focuses on CPU-CPU communication, as it is the most common use case in asynchronous multithreaded applications. However, we recognize that GPU-direct communication is becoming increasingly popular and important for LCI’s future work. There are mature techniques for integrating GPU-Direct RDMA into existing communication libraries, such as MPI and GASNet-EX, so we do not expect it to be difficult for LCI. GPU-initiated communication is a different issue, and we are still looking for potential applications and use cases.

### v2

@@ 5.5. Implementation Note
LCI is implemented as a C++17 library with the CMake build system. It is also available as a Spack package. It has been tested on Infiniband, RoCE, Slingshot-11, and Ethernet networks. It is fully open-source with the NCSA license.
@@ 6.1. Experimental Setup
We evaluate LCI on SDSC Expanse (Strande et al., 2021) and NCSA Delta (Gropp et al., 2023). Table 2 shows their configuration. Expanse uses InfiniBand, which is deployed in 61% of the Top500 systems. Delta uses Slingshot-11, which is increasingly popular and used in 7 of the top 10 systems. [^1] Statistics are based on the TOP500 List published in Nov. 2024. All experiments are conducted at least six times. The figures show the average and standard deviation.
@@ 6.2. Micro-benchmarks
In asynchronous multithreaded applications, message rate and bandwidth are usually more critical than latency due to communication overlapping and nonblocking execution. Therefore, we use these two metrics to compare LCI with standard MPI, MPICH with the VCI extension, and GASNet-EX. Our micro-benchmarks run on two nodes with two basic modes. The process-based mode uses one process on each core, while the thread-based setting uses one process on each node with one thread per core. Each process/thread has a peer process/thread on the other node, and it performs ping-pongs with the peer. Existing multithreaded applications can either share a global set of communication resources or, if the application logic and underlying communication library permit, allocate dedicated resources for each thread. Therefore, the thread-based mode is further divided into two sub-modes according to the resource-sharing pattern: (a) in the dedicated resource mode, each thread allocates its communication resources; (b) in the shared resource mode, all threads share a global set of communication resources. The dedicated resource mode is implemented with MPICH VCIs and LCI devices. Cray-MPICH and GASNet-EX do not support this mode. We also set mpi_assert_no_any_tag and mpi_assert_allow_overtaking to true and configure MPIR_CVAR_CH4_GLOBAL_PROGRESS to 0 to minimize the thread contention on VCIs.
To ensure uniformity across different communication libraries, we build a simple layer (the Lightweight Communication Wrapper, or LCW) on top of LCI, MPI, and GASNet-EX and use it to write the microbenchmarks. The microbenchmarks and the LCW layer are open-sourced [^2] https://github.com/JiakunYan/lcw. LCW implements simple nonblocking active messages and send-receive primitives. For MPI, it uses MPI_Isend/MPI_Irecv for send-receive and MPI_Isend/pre-posted MPI_Irecv for active messages. For GASNet-EX, it uses gex_AM_RequestMedium for active messages and does not support send-receive due to implementation complexity. We show the active message results in the message rate microbenchmark and the send-receive results in the bandwidth microbenchmark.
6.2.1. Single-thread Performance
@@
Figure 3. Process-based message rate micro-benchmark. We use one process per core, each with a single thread, across two nodes.
Figure 3 shows the single-thread message rate results. We fixed the message size to 8 bytes and increased the process number from 1 to 128 per node. Each process/thread runs 100k iterations. We report the unidirectional message rate. LCI achieves performance comparable to that of the other communication libraries. Figures for the single-thread bandwidth results are omitted due to page limit, but the results are similar.
@@ 6.2.2. Multithreaded Performance
Figure 4. Thread-based message rate micro-benchmark. We use one process per node, with one thread per core, across two nodes. Dedicated resources uses one LCI device/MPICH VCI per thread. Shared resources uses one set of resources for the entire process.
Figure 4 shows the multithreaded message rate results. We fixed the message size to 8 bytes and increased the thread number from 1 to 128 per node. LCI achieves significant speedups in multithreaded performance on both platforms (sometimes more than 10x). In particular, multithreaded LCI with dedicated devices achieves even slightly better performance than multi-process LCI (around 15% at full scale). The MPICH VCI extension greatly helps multithreaded performance, but the overall performance is still suboptimal. GASNet-EX shows good multithreaded performance in the shared resource mode, but its lack of resource-replication support reduces its competencies if the application can leverage more resources.
Even though we do not directly evaluate UCX and Libfabric due to the difficulty of bootstrapping and the complexity of their APIs, the MPICH results on Expanse (particularly Figure 4(a)) give hints at their multithreaded performance. UCX is generally faster than libfabric on InfiniBand, but its performance degrades sharply when there are more than 16 threads. Libfabric shows good scaling results with dedicated resources at the cost of absolute performance numbers. LCI achieves the best of both worlds by directly building on the lowest-level public API, libibverbs. UCX does not support Slingshot-11, so its results on Delta are unavailable. MPICH does not support more than 64 VCIs, so some data points are missing.
We have also tested administrator-installed OpenMPI on Expanse. In the shared resource mode with 64 threads, it performs roughly 2x better than MPICH with UCX, but still 5x worse than LCI. It does not support the dedicated resource mode. Due to space constraints, we omit the full results from the figures.
@@
Figure 5. Thread-based bandwidth micro-benchmark. We use one process per node, with one thread per core, across two nodes. Dedicated resources uses one LCI device/MPICH VCI per thread. Shared resources uses one set of resources for the entire process.
@@ 6.2.3. Individual Resources
LCI communications involve operations on a variety of resources. Each resource is optimized for threading efficiency, and users can explicitly allocate multiple replicas of them. Our next microbenchmark evaluates the threading efficiency of three major LCI resources: completion queue, matching engine, and packet pool. All microbenchmarks run on a single node on Delta with different thread numbers. All threads perform 100k of key resource methods in the communication critical path (a pair of completion queue push/pop, matching engine inserts, or packet pool get/put). Figure 6 shows the results. As we can see, the packet pool and matching engine scales well with thread number, achieving 800 Mops (Million operations per second) for the packet pool or 260 Mops for the matching engine, with 128 threads. As a reference, our ping-pong microbenchmark achieves at most 22 Million Messages per second (Figure 4(a)). This means allocating one instance of these two resources per process is sufficient. The completion queue achieves 18 Mops with 128 threads, which means applications aiming for higher throughput may need to allocate multiple completion queues per process. The completion queue throughput is primarily constrained by how fast threads can perform the atomic fetch-and-add operation on a shared variable. Our message rate microbenchmark shown above uses one completion queue per thread.
@@ 6.3. K-mer Counting
In the k-mer counting step, HipMer traverses the dataset twice. The first traversal inserts the k-mers into a two-layer Bloom filter. A Bloom filter is a space-efficient data structure that tests whether an element is in a set with a small false positive rate. The second traversal then consults the Bloom filter and inserts those with more than one occurrence into a HashMap. The HashMap maintains the actual count of the k-mers, while the two-layer Bloom filter reduces the memory footprint of the hashtable by filtering out those occurring only once (which are likely erroneous).
HipMer is written in UPC++ and has only one thread per process. Each k-mer is statically mapped to a process using a hash function. Each process reads part of the dataset and sends the k-mers to the mapped processes via UPC++ RPCs. It further employs an aggregation buffer per target process to reduce communication overhead.
We implement a multithreaded version of the HipMer k-mer counting stage. The new implementation is also based on the RPC abstraction and aggregation, with a libcuckoo hash table(Li et al., 2014) and a hand-written atomic-based Bloom filter. It supports two network backends, LCI and GASNet-EX, primarily leveraging their active message primitives. The LCI backend shares many similarities with the one described in Section 4.2. Compared to the single-threaded implementation, multithreading reduces the number of aggregation targets by a factor of N, where N is the number of threads per process. All threads can serve the incoming RPCs, resulting in improved load balance.
@@
We run the k-mer counting mini-app with the human chr14 dataset (7.75GB). It contains 37 million reads and 1.8 billion k-mers (with k-mer length $k=51$). We run the multithreaded implementation with 2 processes per node to avoid the inter-socket overheads. The aggregation buffer size is set to be 8KB per destination. Due to the reduced destination number, the total aggregation buffer size is always smaller than its HipMer counterpart. All threads run the application logic and periodically progress the network backend (the all-worker setup). This is the best setup for LCI on both platforms. However, when running GASNet-EX on Delta, the all-worker setup results in devastating performance (over 20x worse than LCI). Therefore, we add a dedicated progress setup for GASNet-EX: We use 63 threads for application logic and one thread for network progress. We report the better of the two setups for GASNet-EX (the all-worker setup on Expanse and the dedicated progress setup on Delta).
Figure 7 shows the strong scaling results of the mini-app on Expanse and Delta from 1 node (2 processes/128 cores) to 32 nodes (64 processes/4096 cores). We also include the experimental results of the single-threaded reference implementation to highlight the performance benefits of efficient multithreaded communication over the traditional one-process-per-core model. Our multithreaded implementation outperforms the single-threaded reference implementation by up to 60% on Expanse (8 nodes) and 40% on Delta (4 nodes), at which point the reference implementation suffers from severe load imbalance problems across 1024 (Expanse) or 512 (Delta) processes. In addition, the LCI backend outperforms its GASNet-EX counterpart by 35% on Expanse (16 nodes) and 55% on Delta (16 nodes). Although not shown here, we also tried larger aggregation buffer sizes (up to 64KB), resulting in slightly smaller gaps between GASNet-EX and LCI due to less frequent communication and lower overall performance due to worse load balance.
HipMer evaluation stops at 8 nodes on Expanse and 4 nodes on Delta because UPC++ takes too long to bootstrap for larger process counts. Investigation shows it was stuck inside the slow PMI2 fence operation. It reflects another common challenge of the one-process-per-core running mode on modern HPC systems with high intra-node parallelism.
@@ 6.4. HPX and Octo-Tiger
The Asynchronous Many-Task (AMT) model expresses applications as fine-grained tasks with dependencies, enabling better load balance and communication–computation overlap than bulk-synchronous approaches. HPX (Kaiser et al., 2023) is an established AMT runtime that extends the C++ Standard APIs to distributed computing. (Yan et al., 2023) has integrated a previous C version of LCI (Snir et al., 2023) into it. In this work, we upgrade the LCI support inside HPX to the latest C++ version. HPX primarily uses LCI’s active message primitive for short control messages and its send-receive primitive for bulk data transfers. It issues communications with aggressive asynchrony and concurrency and uses LCI completion queues for synchronization. All HPX threads can issue and progress communications, and they leverage multiple LCI devices to reduce contention.
We evaluate this integration using Octo-Tiger (Marcello et al., 2021), an astrophysics application simulating the evolution of stellar systems based on adaptive octo-trees and fast multipole methods. Octo-Tiger is built on top of HPX for fully asynchronous execution and communication overlapping. We use the ”rotating star” scenarios and report time per simulation step.
@@
Figure 8 presents the results. For comparison, we include HPX’s original MPI backend (mpi) and an enhanced version with the MPICH VCI extensions (mpix(Yan et al., 2025b), both using the MPICH libfabric backend, as prior experiments have shown it delivers better performance. Results reported here use the optimal VCI count for mpix and the optimal device count for lci. We also use replicated request pools for mpix to reduce thread contention on completion polling. On Expanse, LCI outperforms mpi (standard MPI) by 30% and mpix by 10%. On Delta, LCI outperforms mpi by 3x and mpix by 35%. In addition, mpix needs 8 VCIs on both platforms to reach the optimal performance, while lci only needs 1 device on Expanse and 2 devices on Delta. This shows that LCI has better intra-resource threading efficiency compared to MPICH, thanks to its thread-efficient runtime design.
(Daiß et al., 2024) scales Octo-Tiger to 1700+ GPU nodes (around 7000 GPUs/processes) on Perlmutter and achieves 1.7x speedup compared to Cray MPICH at full scale by using LCI. We are unable to conduct similar scaling experiments in this paper due to computational resource limitations, but we expect the latest version of LCI to have similar scalability.
@@ 7. Conclusion and Future Work
LCI is currently under active development. There are several areas for future improvements. We list some of the most important ones below.
Native Network Layer Efficiency: While LCI’s communication library layer is designed for threading efficiency, its performance is currently limited by the threading inefficiencies of the underlying native network layer, particularly libfabric. As discussed earlier, libfabric’s coarse-grained locking makes efficient multithreading difficult. Although LCI mitigates this through techniques such as replicated domains and try-lock wrappers, the resulting performance remains suboptimal. We hope to work with native network layer developers to address these threading bottlenecks. Another potential path is to build LCI directly on top of the lower-level libcxi layer, which is in the process of being open-sourced.
GPU Communication: The existing LCI focuses on CPU-CPU communication, as it is the most common use case in asynchronous multithreaded applications and the backbone for most GPU communication libraries. However, we recognize that GPU-direct communication is becoming increasingly popular and important for LCI’s future work. LCI already supports host-initiated GPU-Direct RDMA. We also have a prototype system for device-initiated LCI operations and are actively looking for potential applications and use cases.
The efficient multithreaded communication and flexible asynchronous primitives provided by LCI open up a significantly larger design space for developing new programming models and algorithms that can fully exploit these capabilities. We believe LCI can serve as a foundation for future systems that push the boundaries of asynchronous and multithreaded computing. We look forward to collaborating with researchers and developers of programming models and applications to explore and realize the full potential of these emerging paradigms.
Acknowledgements.
We thank Benjamin Brock and Aydin Buluç for their assistance in developing the initial multithreaded K-mer Counting Mini-app. We thank Rob Egan, Steven Hofmeyr, Dan Bonachea, Paul H. Hargrove, and Katherine A. Yelick for their assistance in running HipMer, UPC++, and GASNet-EX. We thank Patrick Diehl and Hartmut Kaiser for their assistance in running OctoTiger and HPX. This work used Expanse at San Diego Supercomputer Center (Strande et al., 2021) and Delta at National Center for Supercomputing Applications (Gropp et al., 2023) through allocation CCR130058 and CIS250465 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program (Boerner et al., 2023) supported by U.S. National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296.